4 Pole Mission Distortion

Hello, I have some troubles with my Shruthi 4-pole M
In LP1 and LP2 modes, the sound level is very high and very distorded. It sounds like it’s passed thru a Guv’nor distortion pedal…
Pichenettes took a look at my filter board and it seems looks ok.
I need some help to fix it please…
I am not an electronics and my skills are very limited … However, I have an oscilloscope to make measurements if needed.

Too much resonance? it will cause the signal to clip if the resonance is too high.

Pichenettes is the expert, he designed it.

Does the filter sound normally with the (default) LP4 mode?

If you have a scope, boot the synth, leave the init patch, and post scope pictures of the signals at pin 1 of IC1 ; and at those points:

If you suspect the problem has something to do with resonance, remove the LM13700 (without the LM13700, no resonance).

hello, I have removed the LM13700 and it causes no sound from the Shruthi’ Out .
I’ve replaced it with a new one, and it makes the same overdrive in LP1 and LP2
take a look at the pictures I’ve done and let me know what you think about it

picture 2

picture 3

picture 4

not tested IC 1

Yes, removing LM13700 breaks the audio out, but it does not break the filter. I would have prefered seeing the scope traces without LM13700. It’s not because the audio out is broken that the filter is broken!

Anyway, your traces look normal, so the problem is somewhere else.

Check all the solder points around the LM13700 ; and please show the signal traces at pin 8 of IC2 ; with the filter set to LP4, LP2 and LP1.

here are some new test with pictures
the three first pictures are from IC2 Pin8

LP2

LP4

now from the 4 test points

Test point 2 without LM13700

TP3 without LM13700 ( the line are not dual, this is due to my poor oscilloscope and my old cam )

same as the previous, there were not two traces… too long shooting time I think

LP4 is OK.

Can you confirm that the signals on the 1, 2, 3, 4 pads shown above are identical when you switch the filter from LP4 to LP2? And doesn’t change much when you switch the filter from LP4 through LP1?

If this is the case, the filter itself is working, the problem is in the pole mixing circuit - some nodes are working, some others are not. Check the soldering of the big row with all the resistors (above the 3 logic ICs). Maybe you could post a photo of those. Check the soldering of the 4051.

I don’t know if I can really confirm, in fact it changes a bit as you can see on pictures.

I don’t have any pictures to compare the signal on the 1, 2, 3, 4 with the filter set to LP4 and LP2.

here is a picture from the board Assy where you can eventually see all the resistor, maybe I’ve made a mistake ?


( don’t know how to put a link for an Url here . hummm ) ( edited : thank you